OpenStudy (anonymous):

Help MEDAL and Fan!!! The quilt pattern includes 32 small triangles. Each has a base of 3 in. and a height of 1.5 in. Find the amount of fabric used to make the 32 triangles.

OpenStudy (jdoe0001):

well, how do you find the area of ONE triangle whose base is 3 and height is 1.5?

Directrix (directrix):

Is this the diagram? @Pepper5

OpenStudy (anonymous):

Yes @Directrix

Directrix (directrix):

@Pepper5 The area of one triangle can be found by this formula: A = (1/2) * b * h where b is the length of the base of the triangle and h is the height drawn to that base. A = (1/2) * 3 * 3/2) = 9/4 square inches as the area of ONE triangle. To get the amount of material needed for 32 triangles, multiply 32 by (9/4). You can crank that answer out by hand. Or, copy and paste "32 times (9/4)=" in the Google search box. The Google calculator will pop up with the answer. @Pepper5